MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2007/11/24
OBSERVER: Miyake Paul
CONDITION: strong wind -> slightly cloudy -> strong wind
COMMENT: vacuum sensor broken, change vacuum sensor(No.44200127 -> No.44200114), punming before observation, change exposure time 240sec due to full moon, DARK D3173(240sec) is OK
Compressor gas pressure: 1.9 MPa
Expected Observation Time: START 22:34 END 04:33


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB-000
LMC2068-20714143360
SMC475-47511240
ALERT-000
GRB-000
TOTAL---5153600 (01h00m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
00000000000


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111111111111
Field13Field14Field15Field16
1100


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
100000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
22:00waitingstrong wind
23:00waitingstrong wind
00:00observingslightly cloudy
01:00waitingcloudy
02:00observingslightly cloudy
03:00observingslightly cloudy
04:00waitingstrong wind


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day